OpenStudy (anonymous):

Plz help! An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation?

OpenStudy (anonymous):

Let's simplify step by step. -2(4x-1) - 7

OpenStudy (anonymous):

Distribute =(−2)(4x)+(−2)(−1)+−7 =−8x+2+−7

OpenStudy (anonymous):

First distribute. -4x + 2 - 7 = 5 Then combine like terms. -4x - 5 = 5 Add five to both sides -4x = 10 Divide by four to get X alone X = -2.5 Get it?

OpenStudy (anonymous):

A. The equation can become -2(4x - 1) = -2 by applying the distributive property. B. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. C. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition D. The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (anonymous):

Well than....That's a completely different ball park.

OpenStudy (anonymous):

We could start by adding 7 to both sides, giving us -2(4x - 10 = 12

OpenStudy (anonymous):

Combine like terms = -8x + 2 + - 7 = (-8x) + (2 + - 7) = -8x + - 5

OpenStudy (anonymous):

so that would be addition and that will be B, most likely. That help?

OpenStudy (anonymous):

Answer = -8x - 5

OpenStudy (anonymous):

@Thehulk49 We were trying to answer the wrong question, we don't need to solve it.

OpenStudy (anonymous):

Sorry..

OpenStudy (anonymous):

that's ok :), I did the same thing. :)

OpenStudy (anonymous):

yeah kinda and @Thehulk49 i don't think that is one of the answers

OpenStudy (anonymous):

Lol. Sorry. I guess everyone does make mistakes

OpenStudy (anonymous):

@Calvin3 I should think the answer to your question is B, an answer you get by adding 7 to both sides. @Thehulk49 Yes we do, it's alright.

OpenStudy (anonymous):

Thanks for not being mad..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Gotta go, the bell is going to ring in about 4 seconds. :)

OpenStudy (anonymous):

Bye! :D

OpenStudy (anonymous):

byee :)

OpenStudy (anonymous):

byee

OpenStudy (anonymous):

thx for all your help @Thehulk49